John knows that monthly demand for his product follows a normal distribution with a mean of 2,500 units and a standard deviation of 425 units. Given this, please provide the following answers for John.

a. What is the probability that in a given month demand is less than 3,000 units?

b. What is the probability that in a given month demand is greater than 2,200 units?

c. What is the probability that in a given month demand is between 2,200 and 3,000 units?

d. What is the probability that demand will exceed 5,000 units next month?

e. If John wants to make sure that he meets monthly demand with production output at least 95% of the time. What is the minimum he should produce each month?
Answer #1

a)

µ =    2500              
σ =    425              
left tailed                  
X ≤    3000              
                  
Z =   (X - µ ) / σ =   1.18          
                  
P(X ≤   3000   ) = P(Z ≤   1.18   ) =   0.8803 [ excel formula =normsdist(0.8803)

b)

µ =    2500                          
σ =    425                          
right tailed                              
X ≥   2200                          
                              
Z =   (X - µ ) / σ =   -0.71                      
                              
P(X ≥   2200   ) = P(Z ≥   -0.71   ) =   P ( Z <   0.71   ) =    0.7599 [ excel formula =normsdist(0.71)

c)

µ =    2500                              
σ =    425                              
we need to calculate probability for ,                                  
2200   ≤ X ≤    3000                          
X1 =    2200   ,   X2 =   3000                  
                                  
Z1 =   (X1 - µ ) / σ =   -0.706                          
Z2 =   (X2 - µ ) / σ =   1.176                          
                                  
P (   2200   < X <    3000   ) =    P (    -0.705882353   < Z <    1.176   )
                                  
= P ( Z <    1.176   ) - P ( Z <   -0.706   ) =    0.8803   -    0.2401   =    0.6402

d)

µ =    2500                          
σ =    425                          
right tailed                              
X ≥   5000                          
                              
Z =   (X - µ ) / σ =   5.88                      
                              
P(X ≥   5000   ) = P(Z ≥   5.88   ) =   P ( Z <   -5.88   ) =    0.0000

e)

µ =    2500                  
σ =    425                  
proportion=   0.95                  
                      
Z value at    0.95   =   1.645   (excel formula =NORMSINV(α))      
z=(x-µ)/σ                      
so, X=zσ+µ=   1.645   *   425   +   2500  
X   =   3199.06              
so, he should produce minimum 3199.06 units each month